# MonoForce: Learnable Image-conditioned Physics Engine

### v1

@@ IV-B Sensor Fusion
For different sensor inputs (multi-view RGB cameras and LiDAR), modality-specific encoders are applied to extract their features. Then the sensor-specific features are transformed into a unified BEV representation that preserves both semantic and geometric information. Following the [24] approach the convolution-based BEV encoder is applied next to the stacked BEV features to account for possible local misalignment between different sensors. The task-specific heads are added to predict different terrain properties (for example, its shape as a form of heightmap and friction) as shown in the Figure 13. The LSS [22] model is used as the camera input BEV-encoder, as described in Section III-A and the BEVFusion [24] paper. The VoxelNet [23] model is utilized as the point cloud input BEV-encoder. First, the raw point cloud input is voxelized into a $3D$ voxel grid. The point cloud coordinates are mapped to a discrete grid of binary indices (a value of $1$ denotes that the voxel is occupied). Next, the $3D$-convolution-based LiDAR Encoder processes the voxelized input to extract high-level feature representation. This operation is followed by $\max$-pooling along $z$-axis to reduce the $3D$ feature map to $2D$ BEV space.
@@ IV-D Physics Engine Computational Efficiency
The Figure 15 shows the runtime of the differentiable physics engine depending on the predicted trajectories time horizon for the two solvers implementation methods and different hardware configurations. The Figure 15 (a) shows the runtime on the standard laptop hardware (CPU and GPU), while the Figure 15 (b) shows the runtime on the high-performance GPU. It can be noticed that the runtime grows linearly with the time horizon in all cases and the laptop GPU implementation is around 5 times faster than the CPU one. The Neural ODE solver [45] is slightly more efficient than the auto-differentiation one. For example, on a high-performance GPU (Tesla V100-SXM2-32GB), the prediction of 2048 trajectories ($6~{}[$\sec$]$-long) takes around $1~{}[$\sec$]$. The short runtime of the differentiable physics engine allows for its efficient usage in the end-to-end learning pipeline (Section III-D) and for real-time trajectory shooting in navigation scenarios (Section IV-E).
@@ IV-E Navigation in Unstructured Environments
Figure 17: Autonomous navigation in the forest environment with MonoForce. The robot follows a set of waypoints and at the same time avoids obstacles (trees, bushes, and rocks, etc.). First row (Control setup): MonoForce prediction, based only on the onboard camera, includes terrain shape and a set of trajectories for different control commands sampled through a simplified MPPI technique. The colors of a possible trajectory correspond to the cost (red is the most expensive). The trajectory with green arrows is the one with the smallest total cost, (15). Second row (Qualitative results): Predicted supporting terrain and robot trajectory given camera images. The predictions and frontal camera images correspond to different time moments of the navigation experiment in the forest environment (Figure 16).
Figure 18: MonoForce prediction projected to robot’s camera images. The prediction include supproting terrain $H_{t}$; the colors correspond to the friction coefficient $\mu$ (violet is low, red is high). A set of 64 trajectories for different control commands is visualized in white and black (low and high trajectory cost (13) respectively). The black frames denote cameras field of views.
Thanks to integration of the extereoceptive sensor to Terrain Encoder and GPU parallelization of the predicted trajectories (Section III-B), it is possible to use the MonoForce model for autonomous navigation in unstructured outdoor environments. In the navigation experiments, the robot is given a set of distant waypoints to follow. It is localized using the ICP SLAM method [48]. The MonoForce model runs onboard the robot’s hardware and predicts terrain properties in front of the robot and a set of possible trajectories for different control commands. In case of a tracked robot platform (Figure 12), the controls are the commanded velocities of individual tracks. A sample of predicted trajectories for given controls is visualized in Figure 18. To reach a goal safely, the robot chooses the trajectory with the smallest cost and distance to the next waypoint, Figure 17 Thanks to the MonoForce ability to predict robot-terrain interaction forces for a trajectory pose, it is possible to estimate the cost of each trajectory. We decide to calculate the cost as a values variance of reaction forces acting on the robot along a trajectory $\tau$:
@@
where $\alpha$ and $\beta$ are the hyperparameters for the trajectory and waypoint costs, respectively. With the described navigation approach, the robot is able to autonomously navigate in the forest environment with uneven terrain, avoiding obstacles and following the set of waypoints, Figure 17. The Figure 16 shows the constructed point cloud map of the forest environment during the navigation experiment. It also contains an aerial photo of the forest area with the robot’s trajectory and the waypoints. The robot was able to autonomously navigate in the forest environment and traversed the 260-meter-long path.
@@ V Conclusion
In this work, we have presented MonoForce, an explainable, physics-informed, and end-to-end differentiable model that predicts robot’s trajectory from monocular camera images. As a grey-box model (physics-based and data-driven), it benefits from the end-to-end trainability in different domains while it still retains the determinism of its physics engine and explainability of its estimates of terrain properties. Thanks to the method’s end-to-end differentiability, it is possible to incorporate external sensor measurements in the training pipeline. Additionally, due to the inclusion of the physics information, our model is able not only to generalize better (predict more accurate trajectories than data-driven baselines) but also to provide interpretable intermediate outputs, for example friction and robot-terrain interaction forces. The training process is self-supervised; it only requires monocular camera images, lidar scans and SLAM-reconstructed trajectories. It treats them as rigid obstacles until future data prove otherwise.

### v2

@@ IV-B Sensor Fusion
For different sensor inputs (multi-view RGB cameras and lidar), modality-specific encoders are applied to extract their features. Then the sensor-specific features are transformed into a unified BEV representation that preserves both semantic and geometric information. Following the [22] approach the convolution-based BEV encoder is applied next to the stacked BEV features to account for possible local misalignment between different sensors. The task-specific heads are added to predict different terrain properties (for example, its shape as a form of heightmap and friction) as shown in the Figure 13. The LSS [20] model is used as the camera input BEV-encoder, as described in Section III-A and the BEVFusion [22] paper. The VoxelNet [21] model is utilized as the point cloud input BEV-encoder. First, the raw point cloud input is voxelized into a $3D$ voxel grid. The point cloud coordinates are mapped to a discrete grid of binary indices (a value of $1$ denotes that the voxel is occupied). Next, the $3D$-convolution-based lidar Encoder processes the voxelized input to extract high-level feature representation. This operation is followed by $\max$-pooling along the $z$-axis to reduce the $3D$ feature map to $2D$ BEV space.
@@ IV-D Physics Engine Computational Efficiency
The Figure 15 shows the runtime of the differentiable physics engine depending on the predicted trajectories time horizon for the two solvers’ implementation methods and different hardware configurations. The Figure 15 (a) shows the runtime on the standard laptop hardware (CPU and GPU), while the Figure 15 (b) shows the runtime on the high-performance GPU. It can be noticed that the runtime grows linearly with the time horizon in all cases and the laptop GPU implementation is around 5 times faster than the CPU one. The Neural ODE solver [45] is slightly more efficient than the auto-differentiation one. For example, on a high-performance GPU (Tesla V100-SXM2-32GB), the prediction of 2048 trajectories ($6~{}[$\sec$]$-long) takes around $1~{}[$\sec$]$. The short runtime of the differentiable physics engine allows for its efficient usage in the end-to-end learning pipeline (Section III-D) and for real-time trajectory shooting in navigation scenarios (Section IV-E).
@@ IV-E Navigation in Unstructured Environments
Figure 17: Autonomous navigation in the forest environment with MonoForce. The robot follows a set of waypoints and at the same time avoids obstacles (trees, bushes, rocks, etc.). First row (Control setup): MonoForce prediction, based only on the onboard camera, includes terrain shape and a set of trajectories for different control commands sampled through a simplified MPPI technique. The colors of a possible trajectory correspond to the cost (red is the most expensive). The trajectory with green arrows is the one with the smallest total cost, (15). Second row (Qualitative results): Predicted supporting terrain and robot trajectory given camera images. The predictions and frontal camera images correspond to different time moments of the navigation experiment in the forest environment (Figure 16).
Figure 18: MonoForce prediction projected to robot’s camera images. The prediction includes supporting terrain $H_{t}$; the colors correspond to the friction coefficient $\mu$ (violet is low, red is high). A set of 64 trajectories for different control commands is visualized in white and black (low and high trajectory cost (13) respectively). The black frames denote the camera’s field of view.
Thanks to integration of the extereoceptive sensor to Terrain Encoder and GPU parallelization of the predicted trajectories (Section III-B), it is possible to use the MonoForce model for autonomous navigation in unstructured outdoor environments. In the navigation experiments, the robot is given a set of distant waypoints to follow. It is localized using the ICP SLAM method [48]. The MonoForce model runs onboard the robot’s hardware and predicts terrain properties in front of the robot and a set of possible trajectories for different control commands. In the case of a tracked robot platform (Figure 12), the controls are the commanded velocities of individual tracks. A sample of predicted trajectories for given controls is visualized in Figure 18. To reach a goal safely, the robot chooses the trajectory with the smallest cost and distance to the next waypoint, Figure 17 Thanks to the MonoForce ability to predict robot-terrain interaction forces for a trajectory pose, it is possible to estimate the cost of each trajectory. We decide to calculate the cost as a values variance of reaction forces acting on the robot along a trajectory $\tau$:
@@
where $\alpha$ and $\beta$ are the hyperparameters for the trajectory and waypoint costs, respectively. With the described navigation approach, the robot is able to autonomously navigate in the forest environment with uneven terrain, avoiding obstacles and following the set of waypoints, Figure 17. The Figure 16 shows the constructed point cloud map of the forest environment during the navigation experiment. It also contains an aerial photo of the forest area with the robot’s trajectory and the waypoints. The robot was able to autonomously navigate in the forest environment and traverse the 260-meter-long path.
@@ V Conclusion
In this work, we have presented MonoForce, an explainable, physics-informed, and end-to-end differentiable model that predicts the robot’s trajectory from monocular camera images. As a grey-box model (physics-based and data-driven), it benefits from the end-to-end trainability in different domains while it still retains the determinism of its physics engine and explainability of its estimates of terrain properties. Thanks to the method’s end-to-end differentiability, it is possible to incorporate external sensor measurements in the training pipeline. Additionally, due to the inclusion of the physics information, our model is able not only to generalize better (predict more accurate trajectories than data-driven baselines) but also provide interpretable intermediate outputs, for example friction and robot-terrain interaction forces. The training process is self-supervised; it only requires monocular camera images, lidar scans, and SLAM-reconstructed trajectories. The model learns to recognize non-rigid obstacles similar to those that have been driven over in the reference trajectories while keeping the understanding and prediction capabilities for obstacles it has not encountered yet. It treats them as rigid obstacles until future data prove otherwise.